# Learning Adaptive Parallel Reasoning with Language Models

### v1

@@ 2 Related Work
Parallelizing inference with multiple independent language model calls has emerged as another avenue for inference-time scaling towards improved reasoning performance (Cobbe et al., 2021; Wang et al., 2023). While these ensembling methods effectively enhance performance, their lack of coordination across parallel threads results in redundant computation and suboptimal resource utilization. Recent methods have aimed to address this limitation by proposing fixed parallelizable reasoning structures (Press et al., 2023; Yao et al., 2023; Du et al., 2023; Kim et al., 2024; Ning et al., 2024; Zhang et al., 2024; Zhuge et al., 2024; Teng et al., 2025; Wang et al., 2025). However, these proposed search structures (and, in the case of multi-agent reasoning, the communication protocol) are fixed and hand-designed, limiting their flexibility and scalability. In contrast, our approach leverages reinforcement learning to train language models to adaptively structure search at inference time, dynamically allocating compute between parallel and serial operations. Concurrent to our work, PASTA (Jin et al., 2025) and Hogwild! Inference (Rodionov et al., 2025) also explore parallel reasoning. However, PASTA decomposes tasks into parallel sub-tasks but ultimately merges the complete context from each sub-task back into the main inference trajectory, thus not effectively reducing context usage. Meanwhile, Hogwild! Inference employs parallel worker threads for collaborative reasoning, yet it relies exclusively on prompting without any end-to-end optimization. In contrast, our method uniquely integrates supervised training and reinforcement learning, enabling language models to adaptively and efficiently manage parallel explorations and selectively integrate successful outcomes, maximizing reasoning performance within constrained context windows.
Many parallel inference methods for model serving, e.g., RadixAttention (Zheng et al., 2024), use shared prefixes across language model calls for efficiency reasons. APR builds on these optimizations, as parent-child threading mechanism creates natural opportunities for prefix sharing, which further reduces the computational overhead of APR parallelization.
@@ 3.2 Multi-Threading at Inference Time
3.3 Training Models to Adaptively Parallelize their Search
@@ 5 Conclusions, Limitations, and Future Work
  1. 1.
    Extending to pre-trained language models and general tasks. Currently, our experiments are restricted to non-pretrained LMs on Countdown tasks. Moving forward, we plan to extend our methods to general reasoning tasks with pre-trained LMs, which would validate the approach’s broader applicability.
  2. 2.
    Reducing reliance on supervised training. Our current setup requires bootstrapping by mimicking the symbolic solver. We aim to use strong pre-trained LMs as initial checkpoints and, motivated by DeepSeek R1-Zero (DeepSeek-AI, 2025), explore whether we can bypass the supervised training stage and directly apply reinforcement learning.
  3. 3.
    Better orchestration and communication protocols. As a first step, this work uses fork and join operations for inter-thread communication and thread orchestration. We plan to explore other protocols, such as any-to-any messaging, all-to-all communication patterns, or subscription-based methods.
@@ Acknowledgements
The authors would like to thank Nicholas Tomlin, Boyi Li, Sanjay Subramanian, and Ruiqi Zhong for their valuable discussions. This work was partially supported by compute resources provided through Google’s TPU Research Cloud (TRC) and Nvidia.
@@ A.2 Implementation Details
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
A.3 Symbolic Search Algorithm
@@
A.5 Effect of SoS+ Supervised Training Data
@@ A.6 Effect of Temperature
A.7 Differences Between Sequential Tokens and Wall-Clock Time

### v2

@@ 2 Related Work
Parallelizing inference with multiple independent language model calls has emerged as another avenue for inference-time scaling towards improved reasoning performance (Cobbe et al., 2021; Wang et al., 2023). While these ensembling methods effectively enhance performance, their lack of coordination across parallel threads results in redundant computation and suboptimal resource utilization. Recent methods have aimed to address this limitation by proposing fixed parallelizable reasoning structures (Yao et al., 2023; Du et al., 2023; Kim et al., 2024; Grand et al., 2025; Schroeder et al., 2024; Ning et al., 2024; Zhang et al., 2024; Hua et al., 2024; Zhuge et al., 2024; Teng et al., 2025; Wang et al., 2025) often without learning. However, these proposed search structures (and, in the case of multi-agent reasoning, the communication protocol) are fixed and hand-designed, limiting their flexibility and scalability. In contrast, our approach leverages reinforcement learning to train language models to adaptively structure search at inference time, dynamically allocating compute between parallel and serial operations. Concurrent to our work, PASTA (Jin et al., 2025) and Hogwild! Inference (Rodionov et al., 2025) also explore parallel reasoning. However, PASTA decomposes tasks into parallel sub-tasks but ultimately merges the complete context from each sub-task back into the main inference trajectory, thus not effectively reducing context usage. Meanwhile, Hogwild! Inference employs parallel worker threads for collaborative reasoning, yet it relies exclusively on prompting without any end-to-end optimization. In contrast, our method uniquely integrates supervised training and reinforcement learning, enabling language models to adaptively and efficiently manage parallel explorations and selectively integrate successful outcomes, maximizing reasoning performance within constrained context windows.
Many parallel inference methods have shared prefixes across language model calls, which synergize with modern serving framework to improve serving efficiency via prefix caching and reuse (Zheng et al., 2024). APR builds on these optimizations, as parent-child threading mechanism creates natural opportunities for prefix sharing, which further reduces the computational overhead of APR parallelization.
@@ 3.2 Multi-Threading at Inference Time
3.3 Training Models to Adaptively Parallelize their Reasoning
@@ 5 Conclusions, Limitations, and Future Work
1. Extending to pre-trained language models and general tasks. Currently, our experiments are restricted to non-pretrained LMs on Countdown tasks. Moving forward, we plan to extend our methods to general reasoning tasks with pre-trained LMs, which would validate the approach’s broader applicability.
2. Reducing reliance on supervised training. Our current setup requires bootstrapping by mimicking the symbolic solver. We aim to use strong pre-trained LMs as initial checkpoints and, motivated by DeepSeek R1-Zero (DeepSeek-AI, 2025), explore whether we can bypass the supervised training stage and directly apply reinforcement learning.
3. Better orchestration and communication protocols. As a first step, this work uses fork and join operations for inter-thread communication and thread orchestration. We plan to explore other protocols, such as any-to-any messaging, all-to-all communication patterns, or subscription-based methods.
@@ Acknowledgements
The authors would like to thank Nicholas Tomlin, Boyi Li, Li Erran Li, Sanjay Subramanian, Ruiqi Zhong for their valuable discussions. This work was partially supported by compute resources provided through Google’s TPU Research Cloud (TRC) and Nvidia, and an Ai2 Young Investigator Award.
@@ A.2 Implementation Details
A.3 Additional Results with Larger Models
We evaluated both APR and the SoS+ baseline using a 600M-parameter Llama2 model to study scaling with model size. Figure 8 reports accuracy over average total compute (in thousands of tokens), following §4.1. Both methods improve with increased model capacity, and APR maintains a substantial lead over SoS+ at every compute level.
Figure 8: Accuracy vs. average total compute (in thousands of tokens) for 200M and 600M parameter models. APR consistently outperforms SoS+ and scales more strongly with model size.
A.4 Additional Results with Pretrained Models
To verify that APR’s gains extend beyond models trained from scratch, we fine-tuned a pretrained Qwen2.5 1.5 B model on the same SoS+ and APR demonstration data (approximately 4 k supervised steps). As Table 1 shows, APR significantly outperforms the SoS+ baseline on Qwen, mirroring the trend observed with Llama2. This demonstrates that our parallel reasoning framework is agnostic to model family, size, and pretraining status.
Llama2 200M Qwen2.5 1.5B (pretrained)
SoS+ 57.4 % 57.5 %
APR 83.2 % 80.2 %
Table 1: Accuracy of SoS+ and APR when fine-tuning on a pretrained Qwen2.5 1.5 B model versus a Llama 2 200 M model.
A.5 Extended Context Window Experiments
We evaluated APR and SoS+ on a five-number Countdown variant—whose search space is 40× larger—using context budgets up to 8k tokens. Figure 9 shows that APR continues to improve up to about 6k tokens and outperforms SoS+ beyond 3.5k tokens, achieving gains of 7% and 11% at an 8k-token budget.
Figure 9: Accuracy vs. context window size on the five-number Countdown task for SoS+ and APR.
@@ A.5 Extended Context Window Experiments
3:while $\texttt{state\_deque}\neq\emptyset$ do
4:  $\texttt{current\_state}\leftarrow\texttt{state\_deque}.\texttt{popleft}()$
5:  if $\texttt{current\_state}=\texttt{goal\_state}$ then
@@
8:  if $\textsc{is\_promising}(\texttt{current\_state})$ then
9:   $\texttt{result}\leftarrow\textsc{SoS+}(\texttt{current\_state},\texttt{goal\_state})$
10:   if $\texttt{result}=\texttt{"Goal reached"}$ then
@@
14:   $\texttt{states\_to\_explore}\leftarrow\textsc{SE}(\texttt{current\_state})$
15:   $\texttt{state\_deque}.\texttt{extend}(\texttt{states\_to\_explore})$
@@
2:  state_deque $\leftarrow\textsc{SE}(\texttt{start\_state})$
3:  while state_deque $\neq\emptyset$ do
4:   current_state $\leftarrow\texttt{state\_deque}.\texttt{popleft}()$
@@
8:   states_to_explore $\leftarrow\textsc{SE}(\texttt{current\_state})$
@@
10:     parallel_results $\leftarrow[\,\textsc{APR}(s,\texttt{goal\_state},\texttt{False})\mid s\in\texttt{states\_to\_explore}]$
@@
A.6 Symbolic Search Algorithm
@@
A.8 Effect of SoS+ Supervised Training Data
@@ A.9 Effect of Temperature
A.10 Differences Between Sequential Tokens and Wall-Clock Time